OpenStudy (anonymous):

What is the solution set of {x | x > -5} U {x | x < 5}? all numbers except -5 and 5 the numbers between -5 and 5 the empty set all real numbers

OpenStudy (anonymous):

what is x?

OpenStudy (unklerhaukus):

x is greater than -5, and less than 5

OpenStudy (anonymous):

i think the 2nd one

OpenStudy (anonymous):

This question doesn't makes sense with definition of x for example if \(x \in \mathbb {R}\) then "it is the numbers between -5 and 5" if \(x \in \mathbb {Z}\) or \(x \in \mathbb {W}\) or then "the empty set"